OpenStudy (anonymous):

The tension at which a fishing line snaps is commonly called the line's strength .what minimum strength is needed for a line that is to stop a 19 lb salmon in 4.4in.if the fish is intitially drifting at 9.2ft/s?Assume a constant deceleration

OpenStudy (anonymous):

find the decelleration rate use eqns of constant acc to find it then m*g-T=m*a find T a is decellaration

OpenStudy (anonymous):

start by converting everything to SI. 8.62kg fish .112m of distance traveled 2.8m/s initial velocity 0m/s final velocity velocity(final)^2=velocity(initial)^2 +2a(change in x) velocity(final)^2-velocity(initial)^2=2a(change in x) (velocity(final)^2-velocity(initial)^2)/(2(change in x))=a (0-(2.8)^2)/(2(.112)=a 35m/s^2=a F=ma F = 8.62 * 35m/s^2 F = 301.7 N same thing in given(not SI) units approx 183 ft/lb (I think) staying with SI units is probably your best bet